OpenStudy (anonymous):

What is the center of the circle with equation (x - 1)² + (y + 5)² = 9 Question 18 options: a) (1,5) b) (-1,-5) c) (-1,5) d) (1,-5)

OpenStudy (mathstudent55):

Equation of circle with center (h, k) and radius r. \( (x - h)^2 + (x - k)^2 = r^2 \) \( (x - 1)^2 + (y + 5)^2 = 9 \) \( (x - \color{red}{1})^2 + (y - (\color{red}{-5}))^2 = \color{red}{3}^2 \) Circle with center at (1, -5) and radius 3.